# The Interplay Between Logical Phenomena and the Cognitive System of the Mind

### 08 Jan 2024 v1

@@ 1. Preliminaries and Introduction
To gain a comprehensive understanding of the intricate relationship between the chemical and physical effects of the brain and the experience of consciousness, it is indeed crucial to delve into the process of transforming a non-living chemical substance into a living entity and the subsequent creation of a cognitive system. The transition from non-living matter to a living organism involves a complex series of events, including chemical reactions, self-organization, and the emergence of biological structures. These processes give rise to the fundamental building blocks of life, such as proteins, nucleic acids, and membranes, which form the basis for the intricate machinery of a living system. Within this living entity, the creation of a cognitive system involves the organization and integration of various components, including neural networks, sensory organs, and information processing mechanisms. The precise mechanisms through which consciousness arises from these physical and chemical processes remain a subject of ongoing scientific inquiry and philosophical contemplation. By exploring and unraveling the process of converting non-living chemical substances into living organisms and the subsequent emergence of cognitive systems, we can gain valuable insights into the nature of consciousness and the intricate interplay between the physical, chemical, and cognitive aspects of our existence. This exploration offers a pathway towards a deeper understanding of the profound phenomena underlying consciousness and the complex dynamics that give rise to our subjective experiences. As expounded by Pross (2021)[55], the attainment of a sentient chemical system possessing the capacity to evolve and adapt in order to optimize its interaction with the environment assumes paramount importance for the augmentation of persistence. This protracted process, unfolding over a vast expanse of time, ultimately engendered the advent of the bacterial cell, signifying a momentous milestone in the evolutionary trajectory towards more enduring manifestations.
@@
2. The precise origin of the logical structure of a system
@@
These sets provide a framework for understanding the different aspects and components involved in the functioning of the human mind as proposed by Azar.
@@
As the process continues, a new generation of sensitizers are generated which are sensitive to the performance and changes of $X_{2}$. An example of these sensitizers can be represented by a category denoted as $X_{3}$. It is important to note that all of these sensitivities are molecular, chemical, and physical in nature, occurring within $\gamma$. The elements of $X_{3}$ exhibit functional reactivity that depends on the changes of $X_{2}$ elements. It is worth noting that in the process of $\gamma$ evolution, other components such as $X_{3},X_{4},X_{5}$, etc., may arise in parallel as subsequent sensors within $\gamma$. Consequently, the system components, including the sets $X_{1},\cdots,X_{10}$, develop a unique sensitivity to their constituent elements, which is refined over time. As the process continues, a one-to-one correspondence is established between the system components. For instance, if we consider $\gamma$ to be the brain, different frequencies of light received by $X_{1}$ evoke varying and corresponding sensitivities in $X_{2}$, which we perceive (or experience) as different colors. At this juncture, the chemical composition $\gamma$ may appear to be a simple system, and we refer to it as such. The various sensors that give rise to the system’s components ($X_{i}$, where $1\leq i\leq 10$) are collectively referred to as ”advanced sensors of all generations” ($ASAG$), or simply ”sensors” according to Azar paper (2023)[4]. It is important to note that these sensitivities develop incrementally due to a combination of environmental factors external to the system and chemical and physical reactions within the system.
@@
The root of the experience of consciousness is essentially the result of the function of $ASAG$ and $f(\gamma)$, which is described in detail in Azar (2023)[4]. It is important to note that the interface between a system and its environment is facilitated by sensors. These sensors enable the system to establish a correspondence between its experiences and the changes and events that occur in nature. In this process, the experience of a system is dependent on the surrounding changes and the behavior of its sensors. It is crucial to evaluate the reality of a system’s experience within its own context, as the perception of reality can vary between different systems due to their diverse sensors and components. For instance, most animals and insects are incapable of perceiving the roundness of the Earth or its movement around the Sun because their sensors lack the necessary capabilities. Similarly, humans face similar limitations in comprehending certain phenomena. Our experience of light frequencies, for example, manifests as the observation of different colors. This phenomenon is directly related to the reactions of sensors in components such as $X_{1}$ and $X_{2}$. If our sensors operated differently, our experience of perceiving light would be altered accordingly. This principle holds true for other types of experiences as well. In summary, the interface between a system and its environment is established through sensory mechanisms. The system’s experiences are influenced by the surrounding changes and the behavior of its sensors. The perception of reality must be evaluated within the context of each system, as different systems possess varying sensors and components. This leads to different interpretations and understandings of the surrounding world. Our experience of phenomena, including light frequencies, is shaped by the reactions of specific sensors.
3. On the Mathematical methods for studying of the consciousness experience
Philosophers frequently contemplate the question of whether the experience of consciousness can be precisely defined. If we were to entertain the notion that a comprehensive definition of consciousness is within reach, it would inherently be an experience of consciousness in and of itself. Moreover, if such a definition were to be attained, it would need to possess the ability to encompass and express the entirety of our experiences through its inherent characteristics. Considering that this definition encompasses all phenomena, it implies that the following phenomena can be categorized as experiences. Denoting the set as $X=\{A\notin A:A~{}\text{is a set}\}$, it is evident that $\{1\}$ and $\{1,2\}$ belong to $X$, indicating that $X$ is not empty. Now, regarding $X$ as an experience, we raise the question of whether $X\in X$ or $X\notin X$. Logically, this question should have an answer, but none can be found. Through mathematical analysis, I demonstrate that it is not possible to define all of our experiences with specific characteristics.
@@ 3. On the Mathematical methods for studying of the consciousness experience
In the following, we demonstrate that the occurrence of all physical signals in the brain is unbounded. It is known that for a conscious experience in the mind, there is at least one physical signal that occurs in the brain. Let $\mathcal{P}$ represent the set of all physical signals occurring in the brain, and $\mathcal{E}$ represent the set of conscious experiences in the mind. Thus, there exists a function $\Gamma$ from $\mathcal{P}$ onto $\mathcal{E}$ such that for any conscious experience $Y\in\mathcal{E}$, there exists $X\in\mathcal{P}$ satisfying $\Gamma(X)=Y$. It is evident that the set of conscious experiences in the mind is unbounded. For instance, for each natural number $n$, there exists a conscious experience denoted as $\alpha_{n}$, and there exists a relationship $n\longleftrightarrow\alpha_{n}$ between them. Since the set of natural numbers is infinite, the set of conscious experiences derived from numbers is also infinite. Thus, the occurrence of all physical signals in the brain is unbounded. This outcome indicates that the information stored in our memory is not constrained. Consequently, the processing and analysis of information in our minds will not be limited.
@@
The collection of all experiences of a complete system cannot be a set, as it would lead to a contradiction in set theory. Assuming that $\mathcal{E}$ is the set of all experiences of a complete system, we would have $\mathcal{E}\in\mathcal{E}$, which contradicts the axioms of set theory. To address this issue, we can consider an alternative approach for the collection of system experiences. If we view all humans as a system $w$, we know that each individual human has a finite number of experiences within a given time interval $[t_{1},t_{2}]$. If we define $\mathcal{E}(t)$ as the collection of all experiences of $w$ up to time $t$, then $\mathcal{E}(t)$ is finite. It becomes evident that $\mathcal{E}(t_{1})\subseteq\mathcal{E}(t_{2})$ whenever $t_{1}\leq t_{2}$. We can then define $\mathcal{E}$ as the union of all $\mathcal{E}(t)$, denoted as $\mathcal{E}=\bigcup_{t}\mathcal{E}(t)$. This implies that $\mathcal{E}$ is a countable set representing the possible experiences of the system. By adopting this approach, we avoid the contradiction arising from considering the collection of all experiences as a single set and instead recognize that experiences can be organized and categorized within a temporal framework, leading to a countable set of potential experiences for the system.
@@
The approach you’re describing involves defining a hierarchy of sets $A_{0},A_{1},A_{2},\ldots$ to represent different levels of experiences. The set $A_{0}$ consists of experiences that do not result from other experiences, such as seeing colors, feeling pain, or hearing sounds. The set $A_{1}$ represents experiences that result from the experiences in $A_{0}$, and $A_{n}$ result from the experiences in $A_{0},A_{1},\ldots,A_{n-1}$. However, when we define $\mathcal{E}=\cup_{i=1}^{+\infty}A_{i}$, we encounter a difficulty in treating $\mathcal{E}$ as a set. This is because the process of taking the union of infinitely many sets may not necessarily result in a set according to the axioms of set theory. This issue is related to the concept of a ”set of all sets,” which leads to paradoxes in set theory, such as Russell’s paradox. The paradox arises when considering whether a set can contain itself as an element. In our case, $\mathcal{E}$ would contain all sets $A_{n}$ as its elements, including itself, which leads to a contradiction. To avoid this paradox, one possible approach is to consider the hierarchy of sets $A_{0},A_{1},A_{2},\ldots$ as a potential progression of experiences rather than a set-theoretic construction. In this view, $\mathcal{E}$ would represent the collection of all possible experiences, but it would not be treated as a set in the traditional sense. Instead, it would be an abstract concept denoting the entirety of possible experiences within the system. By reframing $\mathcal{E}$ as a concept rather than a set, we can still discuss the progression of experiences and their relationships without running into the paradoxes associated with the set of all sets.
@@
Indeed, as we navigate through the realm of our experiences, it becomes increasingly apparent that delineating a clear boundary between darkness and light proves to be a formidable task. In simple terms, when we observe the gradual shift from luminosity to obscurity or vice versa, our perceptions do not readily lend themselves to establishing a definitive moment of demarcation. We find ourselves unable to pinpoint an exact point in time where we can confidently assert that prior to a specific moment ”t”, there existed light, and subsequent to that moment, darkness ensued. The fluidity of this transition defies our attempts to impose rigid divisions within the continuum of our perceptual journey.
@@
Let us consider that $\mathcal{E}$ itself represents an experience. In this context, we can define the set $\mathcal{H}$ as the collection of times denoted by $t$ when $\mathcal{E}$ does not belong to the set of experiences at time $t$, represented as $\mathcal{E}(t)$. Mathematically, we can express this as:
@@
It is evident that $\mathcal{H}$ is a non-empty set. Since $\mathcal{H}$ is bounded above, we can establish its supremum as $t_{0}$. Consequently, for any $t>t_{0}$, $\mathcal{E}$ belongs to $\mathcal{E}(t)$. Furthermore, the creation of $\mathcal{E}$ arises from the collection of all our experiences at time $t_{0}$, denoted as $\mathcal{E}(t_{0})$. Assuming that $\mathcal{E}$ can analyze $\mathcal{E}$, $\mathcal{E}(t)$ can produce $\mathcal{E}$ for all $t>t_{0}$. However, $\mathcal{E}(s)$ cannot produce $\mathcal{E}$ for all $s<t_{0}$. Now, let us consider $s<t_{0}<t$. When examining the time interval $[s,t]$, we notice that our experiences will be severely limited when $t$ and $s$ are extremely close together. This leads to a contradiction in our assumption.
4. How are logical and conceptual phenomena extracted from experiential phenomena?

### 21 Jan 2024 v2

@@ 1. Preliminaries and Introduction
To gain a comprehensive understanding of the intricate relationship between the chemical and physical effects of the brain and the experience of consciousness, it is indeed crucial to delve into the process of transforming a non-living chemical substance into a living entity and the subsequent creation of a cognitive system, see Azar (2023)[4]. The transition from non-living matter to a living organism involves a complex series of events, including chemical reactions, self-organization, and the emergence of biological structures. These processes give rise to the fundamental building blocks of life, such as proteins, nucleic acids, and membranes, which form the basis for the intricate machinery of a living system. Within this living entity, the creation of a cognitive system involves the organization and integration of various components, including neural networks, sensory organs, and information processing mechanisms. The precise mechanisms through which consciousness arises from these physical and chemical processes remain a subject of ongoing scientific inquiry and philosophical contemplation. By exploring and unraveling the process of converting non-living chemical substances into living organisms and the subsequent emergence of cognitive systems, we can gain valuable insights into the nature of consciousness and the intricate interplay between the physical, chemical, and cognitive aspects of our existence. This exploration offers a pathway towards a deeper understanding of the profound phenomena underlying consciousness and the complex dynamics that give rise to our subjective experiences. As expounded by Pross (2021)[60], the attainment of a sentient chemical system possessing the capacity to evolve and adapt in order to optimize its interaction with the environment assumes paramount importance for the augmentation of persistence. This protracted process, unfolding over a vast expanse of time, ultimately engendered the advent of the bacterial cell, signifying a momentous milestone in the evolutionary trajectory towards more enduring manifestations. According to Pascal and Pross (2023)[59], even bacteria, which represent the simplest form of prokaryotic life, possess the ability to sense and respond to a wide range of environmental signals. These organisms can sense their internal conditions, coordinate with neighboring organisms, and activate complex response systems to address ongoing challenges, see Shapiro (2007)[63].
Cognitive systems emerge through the interaction of chemical sensors within certain compounds with the external environment and internal processes. As this process continues, it can give rise to unique characteristics in certain chemical compounds, resulting in the formation of living organisms. The ongoing development of this process can lead to the emergence of organisms with more advanced cognitive mechanisms, which we commonly refer to as the mind. The mind encompasses a range of cognitive abilities such as perception, memory, learning, and decision-making. These cognitive capabilities are closely linked to the biological structure of organisms, including the development of complex nervous systems and brain structures that support higher-order cognitive functions. It is through the interplay of these cognitive mechanisms and the biological structures that organisms are capable of exhibiting more sophisticated behaviors, adapting to their environment, and engaging in complex cognitive processes.
Indeed, the experience of consciousness can be seen as an evolutionary development of the cognitive mechanism in living organisms. This process began with simple living systems and may have even emerged from non-living chemical systems. Over time, this cognitive mechanism has evolved and become more complex, culminating in the intricate consciousness experienced by humans and other advanced organisms. Consciousness is believed to have emerged gradually in the course of biological evolution. It likely began in simple living systems with rudimentary forms of perception and awareness. Even single-celled organisms exhibit basic sensory responses and the ability to navigate their environment, suggesting the presence of some level of consciousness. As life forms evolved and became more complex, so did their cognitive abilities. The development of nervous systems and the emergence of more sophisticated sensory organs allowed for enhanced information processing and perception. This evolutionary trajectory led to the gradual refinement and expansion of conscious experiences.
@@
2. Exploring the Root of Conscious Experience
@@
According that manuscript, these sets provide a framework for understanding the different aspects and components involved in the functioning of the human mind.
@@
As the process continues, a new generation of sensitizers are generated which are sensitive to the performance and changes of $X_{2}$. An example of these sensitizers can be represented by a category denoted as $X_{3}$. It is important to note that all of these sensitivities are molecular, chemical, and physical in nature, occurring within $\gamma$. The elements of $X_{3}$ exhibit functional reactivity that depends on the changes of $X_{2}$ elements. It is worth noting that in the process of $\gamma$ evolution, other components such as $X_{3},X_{4},X_{5}$, etc., may arise in parallel as subsequent sensors within $\gamma$. Consequently, the system components, including the sets $X_{1},\cdots,X_{10}$, develop a unique sensitivity to their constituent elements, which is refined over time. As the process continues, a one-to-one correspondence is established between the system components. For instance, if we consider $\gamma$ to be the brain, different frequencies of light received by $X_{1}$ evoke varying and corresponding sensitivities in $X_{2}$, which we perceive (or experience) as different colors. At this juncture, the chemical composition $\gamma$ may appear to be a simple system, and we refer to it as such. The various sensors that give rise to the system’s components ($X_{i}$, where $1\leq i\leq 10$) are collectively referred to as ”advanced sensors of all generations” ($ASAG$), or simply ”sensors” according to Azar’s paper (2023)[5]. It is important to note that these sensitivities develop incrementally due to a combination of environmental factors external to the system and chemical and physical reactions within the system.
@@
The root of the experience of consciousness is essentially the result of the function of $ASAG$ which is described in detail in Azar (2023)[5]. It is important to note that the interface between a system and its environment is facilitated by sensors. These sensors enable the system to establish a correspondence between its experiences and the changes and events that occur in nature. In this process, the experience of a system is dependent on the surrounding changes and the behavior of its sensors. It is crucial to evaluate the reality of a system’s experience within its own context, as the perception of reality can vary between different systems due to their diverse sensors and components. For instance, most animals and insects are incapable of perceiving the roundness of the Earth or its movement around the Sun because their sensors lack the necessary capabilities. Similarly, humans face similar limitations in comprehending certain phenomena. Our experience of light frequencies, for example, manifests as the observation of different colors. This phenomenon is directly related to the reactions of sensors in components such as $X_{1}$ and $X_{2}$. If our sensors operated differently, our experience of perceiving light would be altered accordingly. This principle holds true for other types of experiences as well. In summary, the interface between a system and its environment is established through sensory mechanisms. The system’s experiences are influenced by the surrounding changes and the behavior of its sensors. The perception of reality must be evaluated within the context of each system, as different systems possess varying sensors and components. This leads to different interpretations and understandings of the surrounding world. Our experience of phenomena, including light frequencies, is shaped by the reactions of specific sensors.
In the evolution of the cognitive mechanism, there is no clear demarcation point that separates the transition from non-living to living organisms (Azar (2023)[4]) or from a state without consciousness to a state with consciousness (Azar (2023)[5, 6]). Instead, it is generally considered to be a gradual and continuous process. The evolution of consciousness is believed to have occurred gradually over an extended period. It is not attributed to a single defining moment or specific genetic event but rather a cumulative result of biological and cognitive changes over time. Simple organisms possessed basic sensory and reactive capacities, which gradually developed into more sophisticated cognitive abilities in complex organisms.
The fundamental elements of sensory experiences are qualia, which refer to qualitative sensations or sense data. Qualia are considered the smallest units of phenomenal experience. The mind utilizes these units to form a detailed representation of the external world. The accuracy of these representations determines the extent to which they correspond to the actual external world, thereby constituting veridical perceptions. These states acquire their syntactic properties by virtue of their manifestation in the brain. In other words, qualia, or qualitative sensations, refer to the subjective aspects of our sensory experiences. They encompass the raw, immediate qualities of perception, such as the redness of an apple, the sweetness of sugar, or the warmth of sunlight on our skin. Qualia are highly personal and subjective, as they are experienced individually and cannot be directly shared or communicated with others. They are considered the basic building blocks of our conscious experiences. In every scientific discipline, a consistent set of subject axioms and methods of reasoning is necessary to address the questions and challenges within that field. If we consider qualia as the foundational axioms for the set of conscious experiences, then the definition of these axioms requires additional fundamental axioms that maintain consistency within the system. When analyzing qualia, we rely on the conscious experiences of the system, which are derived from qualia itself. As a result, certain questions related to the thesis, such as defining the experience of consciousness within philosophical frameworks, become difficult to answer in isolation. It becomes necessary to integrate insights from other scientific disciplines such as biology, mathematics, physics, or chemistry. By combining knowledge from these diverse fields, we can gain a more comprehensive understanding of the nature of consciousness and develop a systematic analysis that takes into account various perspectives and empirical evidence. This interdisciplinary approach is crucial to tackle the complex questions posed by qualia and to ensure a more holistic understanding of conscious experiences.
3. Mathematical Approaches in the Study of Consciousness Experience
@@
The question of defining consciousness and understanding its nature has been a subject of deep contemplation and debate among philosophers, neuroscientists, and cognitive scientists. While there is no universally agreed-upon definition of consciousness, it generally refers to the subjective experience of being aware of oneself and the surrounding world. Philosophers often ponder the question of whether the concept of consciousness can be precisely defined. If we entertain the idea that a comprehensive definition of consciousness is attainable, it would inherently be an experience of consciousness in itself. Furthermore, if such a definition were to be achieved, it would need to be capable of encompassing and expressing the entirety of our experiences through its inherent characteristics.
One of the challenges in defining consciousness precisely is that it is an experiential phenomenon. Any attempt to define consciousness would inherently involve experiencing it, which creates a circularity in the definition. In other words, in order to define consciousness comprehensively, one would need to have a conscious experience of that definition. Moreover, if we assume that a comprehensive definition of consciousness is achievable, it would need to encompass and express the entirety of our subjective experiences. Consciousness is not limited to a single aspect or isolated phenomenon but encompasses a wide range of mental states, perceptions, thoughts, emotions, and sensory experiences. Any definition that aims to capture the full scope of consciousness would need to account for these diverse aspects.
When considering this all-encompassing definition, it implies that various phenomena can be categorized as experiences. Let’s denote the set of such phenomena as $X=\{A\notin A:A~{}\text{is a set}\}$. It becomes evident that $\{1\}$ and $\{1,2\}$ belong to $X$, suggesting that $X$ is not empty. Now, viewing $X$ as an experience, we raise the question of whether $X\in X$ or $X\notin X$. Logically, this question should have an answer, yet none can be found. Through mathematical analysis, I demonstrate that it is not possible to define all of our experiences with specific characteristics. Consciousness phenomena are not finite or quantifiable in terms of numbers. Instead, the transition from one phenomenon to another is a continuous process, demonstrating a cardinality comparable to that of real numbers.
In the following, we demonstrate that the occurrence of all physical signals in the brain is unbounded. It is known that for a conscious experience in the mind, there is at least one physical signal that occurs in the brain. Let $\mathcal{P}$ represent the set of all physical signals occurring in the brain, and $\mathcal{E}$ represent the set of conscious experiences in the mind. Thus, there exists a function $\Gamma$ from $\mathcal{P}$ onto $\mathcal{E}$ such that for any conscious experience $Y\in\mathcal{E}$, there exists $X\in\mathcal{P}$ satisfying $\Gamma(X)=Y$. It is evident that the set of conscious experiences in the mind is unbounded. For instance, for each natural number $n$, there exists a conscious experience denoted as $f(n)$, and there exists a relationship $n\longleftrightarrow f(n)$ between them. Since the set of natural numbers is infinite, the set of conscious experiences derived from numbers is also infinite. Thus, the occurrence of all physical signals in the brain is unbounded. This outcome indicates that the information stored in our memory is not constrained. Consequently, the processing and analysis of information in our minds will not be limited.
@@
The collection of all experiences of a complete system cannot be a set, as it would lead to a contradiction in set theory. Assuming that $\mathcal{E}$ is the set of all experiences of a complete system, we would have $\mathcal{E}\in\mathcal{E}$, which contradicts the axioms of set theory. To address this issue, we can consider an alternative approach for the collection of system experiences. If we view all humans as a system, we know that each individual human has a finite number of experiences within a given time interval $[t_{1},t_{2}]$. If we define $\mathcal{E}(t)$ as the collection of all experiences up to time $t$, then $\mathcal{E}(t)$ is finite. It becomes evident that $\mathcal{E}(t_{1})\subseteq\mathcal{E}(t_{2})$ whenever $t_{1}\leq t_{2}$. We can then define $\mathcal{E}$ as the union of all $\mathcal{E}(t)$, denoted as $\mathcal{E}=\bigcup_{t}\mathcal{E}(t)$. This implies that $\mathcal{E}$ is a countable set representing the possible experiences of the system. By adopting this approach, we avoid the contradiction arising from considering the collection of all experiences as a single set and instead recognize that experiences can be organized and categorized within a temporal framework, leading to a countable set of potential experiences for the system.
@@
The approach you’re describing involves defining a hierarchy of sets $A_{0},A_{1},A_{2},\ldots$ to represent different levels of experiences. The set $A_{0}$ consists of experiences that do not result from other experiences, such as seeing colors, feeling pain, or hearing sounds which are called qualia or subjective experiences. The set $A_{1}$ represents experiences that result from the experiences in $A_{0}$, and $A_{n}$ result from the experiences in $A_{0},A_{1},\ldots,A_{n-1}$. However, when we define $\mathcal{E}=\cup_{i=1}^{+\infty}A_{i}$, we encounter a difficulty in treating $\mathcal{E}$ as a set. This is because the process of taking the union of infinitely many sets may not necessarily result in a set according to the axioms of set theory. This issue is related to the concept of a ”set of all sets,” which leads to paradoxes in set theory, such as Russell’s paradox. The paradox arises when considering whether a set can contain itself as an element. In our case, $\mathcal{E}$ would contain all sets $A_{n}$ as its elements, including itself, which leads to a contradiction. To avoid this paradox, one possible approach is to consider the hierarchy of sets $A_{0},A_{1},A_{2},\ldots$ as a potential progression of experiences rather than a set-theoretic construction. In this view, $\mathcal{E}$ would represent the collection of all possible experiences, but it would not be treated as a set in the traditional sense. Instead, it would be an abstract concept denoting the entirety of possible experiences within the system. By reframing $\mathcal{E}$ as a concept rather than a set, we can still discuss the progression of experiences and their relationships without running into the paradoxes associated with the set of all sets.
@@
Indeed, as we navigate through the realm of our experiences, it becomes increasingly apparent that delineating a clear boundary between darkness and light proves to be a formidable task. In simple terms, when we observe the gradual shift from luminosity to obscurity or vice versa, our perceptions do not readily lend themselves to establishing a definitive moment of demarcation. We find ourselves unable to pinpoint an exact point in time where we can confidently assert that prior to a specific moment ”$t$” there existed light, and subsequent to that moment, darkness ensued. The fluidity of this transition defies our attempts to impose rigid divisions within the continuum of our perceptual journey.
@@
Let’s consider $\mathcal{E}$ to represent an individual experience. In this context, we can define the set $\mathcal{H}$ as the collection of times denoted by $t$ when $\mathcal{E}$ does not occur, which we represent as $\mathcal{E}(t)$. Mathematically, we can express this as follows:
@@
It is clear that $\mathcal{H}$ is a non-empty set. Since $\mathcal{H}$ is bounded above, we can establish its supremum as $t_{0}$. Consequently, for any $t>t_{0}$, $\mathcal{E}$ belongs to $\mathcal{E}(t)$. Furthermore, the creation of $\mathcal{E}$ arises from the collection of all our experiences at time $t_{0}$, denoted as $\mathcal{E}(t_{0})$. Assuming that $\mathcal{E}$ can analyze $\mathcal{E}$, $\mathcal{E}(t)$ can produce $\mathcal{E}$ for all $t>t_{0}$. However, $\mathcal{E}(s)$ cannot produce $\mathcal{E}$ for all $s<t_{0}$.
Now, let’s consider $s<t_{0}<t$. When examining the time interval $[s,t]$, we notice that our experiences will be severely limited when $t$ and $s$ are extremely close to each other. This leads to a contradiction in our assumption.
4. Extracting Logical and Conceptual Phenomena from Experiential Realms
Living organisms are formed through an ongoing process of chemical and physical interactions involving specific chemical compounds. These compounds have evolved within cognitive systems, which possess distinct characteristics within the system of living organisms. This information is supported by the article ”Azar (2023)[4]” and its associated sources.
Cognitive systems undergo a transformative process, transitioning between chemical and biological states, encompassing both non-living and living entities. These systems consist of a fusion of chemical sensors, $ASAG$ (Azar (2023)[5, 6]), which emerge through the interplay of assessing the chemical composition of the external environment and internal chemical and physical processes. As a chemical compound evolves into a living organism, the cognitive mechanism responsible for this transformation is further amplified and refined. According to Pamela et al. (2021)[52], basal cognition refers to the essential processes and mechanisms that enabled organisms to monitor specific environmental conditions and respond adaptively to ensure their survival (e.g., finding food, avoiding danger) and reproductive success long before the emergence of complex nervous systems, let alone central nervous systems. Basal cognition implies an inherent level of implicit familiarity or comprehension of the connections between environmental states. In the pursuit of enhanced thermodynamic stability, which forms the foundation for all physical and chemical processes, one of the primary avenues is through the well-established principles of thermodynamics. Azar (2023)[4] posits that distinguishing between living and non-living organisms based on their characteristics is not feasible. The transition from a non-living chemical compound to a living entity is a continuous process without a clear demarcation point. In their manuscript (Pascal and Pross, 2023)[53], the authors propose that perception arises from the dynamic interplay between the replicative system known as the Dynamic Kinetic Stability (DKS) system and its supportive environment. The DKS system is considered a kinetic phenomenon associated with the self-organization of life, driven by the power of replication. Pascal and Pross argue that the DKS system, through kinetic selection, ”learns” which variations are advantageous and which are not. This capacity for replication allows the system to both remember and learn, indicating a mental dimension. Consequently, the concept of DKS provides a framework for understanding the physico-chemical basis of the mental aspects of life.